OpenStudy (anonymous):

Can anyone help with a geometry question? (Questions is inside) [WILL Medal & Fan]

OpenStudy (anonymous):

@Mehek14 quick help?

OpenStudy (anonymous):

is it multiple choice

OpenStudy (anonymous):

The photo of it is right there @Emmily1127

OpenStudy (anonymous):

i saw it

OpenStudy (anonymous):

its the third one

OpenStudy (anonymous):

How'd you get that? What made you choose the third one?

OpenStudy (anonymous):

because on one side of the square there is a 2 so there would be one on the other side and the r stands for the dot so its the 2 2s added equals four plus the dot

OpenStudy (anonymous):

Ok, I just don't want to get it wrong. Can you help with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

im not quite sure but i think its 84

OpenStudy (anonymous):

@phi have an idea?

OpenStudy (phi):

I don't understand your first question, the area of the "shaded sector of the circle" The only "shade" I see is the blue and that is not part of the circle. The area of the blue is the area of the square (= 2*2 or 4) minus the area of of the circle the circle has a radius of 1 (½ of the square's side), and so its area is pi*r*r = pi*1*1= pi so the area of the blue is 4-pi

OpenStudy (phi):

for the two circles in a rectangle, they say the area of each circle is 196 pi the formula for the area of a circle is pi r^2 so write \[ \pi r^2 = 196 \pi\] divide both sides by pi to get \[ \frac{\pi}{\pi}r^2= 196\frac{\pi}{\pi} \\ r^2=196\] you can find the radius of the circle by take the square root of both sides

OpenStudy (anonymous):

So the first one wouldn't be C, it'd be B.

OpenStudy (phi):

Probably, but I would complain to your teacher because they are asking for the area of the sector of a circle, and the blue is not the area of the circle (but I think that is what they want)

OpenStudy (phi):

can you find r using \[ r^2= 196\]?

OpenStudy (anonymous):

Uh yeah, didn't you say find the square root of 196? That would be 14

OpenStudy (phi):

yes, and r is ½ of the diameter. so the diameter of the circle is 2*14 = 28 now we can find the sides of the rectangle it is one circle wide so 28 wide and two circles tall so 2*28 = 56 the perimeter is the sum of the distance all the way round, so 28+56 +28+56

OpenStudy (anonymous):

168 Units.. Which is D!